# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 5. $(e)$
Since we assumed that $[H:H_{0}]=2$, the group $\mathrm{Gal}(H/H_{0})$ is cyclic of order two. The group $\mathrm{Gal}(H_{0}/{\mathbb{Q}})$ is isomorphic to the image of $\varepsilon_{f}$ (which contains $-1$), so it is a cyclic group with even cardinality $n$, and $\mathrm{Gal}(H/{\mathbb{Q}})$ thus has cardinality $2n$. Fix $\sigma\in\mathrm{Gal}(H/{\mathbb{Q}})$ such that $\varepsilon_{f}(\sigma)$ has order $n$. Then $\sigma^{n}\in\mathrm{Gal}(H/H_{0})$, and $\mathcal{B}(\sigma^{n},\cdot)=\mathcal{B}(\sigma,\cdot)^{n}=1$, so by the above, $\sigma^{n}$ is trivial. Thus $\sigma_{|H_{0}}\mapsto\sigma$ defines a splitting of the above exact sequence, and therefore $\mathrm{Gal}(H/{\mathbb{Q}})={\mathbb{Z}}/2{\mathbb{Z}}\sigma_{0}\oplus{\mathbb{Z}}/n{\mathbb{Z}}\sigma$. Hence, representatives for $\mathrm{Gal}(H/{\mathbb{Q}})/M$ are given (possibly with repetitions) by $\sigma_{0},\sigma_{0}\sigma,\sigma,\mathrm{id}$. If $\sigma$ (or $\sigma\sigma_{0}$) is in $M$, the previous paragraph shows that $\varepsilon_{f}(\sigma)=\varepsilon_{f}(\sigma\sigma_{0})\in L^{\times 2}$. Since $\varepsilon_{f}(\sigma)$ is a generator of the image of $\varepsilon_{f}$, this implies that $\varepsilon_{f}$ takes its values in $L^{\times 2}$. Hence, in the case $(c)$, $M$ does not contain $\sigma$ or $\sigma_{0}\sigma$, hence $|B|>2$ and we are done.
@@
When $(e)$ holds, one has $\sigma_{0}c\notin M$, so, since $\sigma_{0}\notin M$, $\sigma_{0}$ is not in the subgroup generated by $M$ and $c$, so the conclusion follows. Furthermore, if $(d)$ holds, then $c\in M$, so $\sigma_{0}c\notin M$, so $(e)$ holds, i.e. $(d)$ implies $(e)$.
@@
If $i$ is not a value taken by $\varepsilon_{f}$ (which is for instance the case if $\varepsilon_{f}^{2}=1$), but $i\in L$, then $n$ is not divisible by $4$. Because $\mathrm{Gal}(H/{\mathbb{Q}})$ is isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}\oplus{\mathbb{Z}}/n{\mathbb{Z}}$, the $2$-torsion subgroup of $\mathrm{Gal}(H/{\mathbb{Q}})$ is isomorphic to ${\mathbb{F}}_{2}^{\oplus 2}$ and projects isomorphically on $\mathrm{Gal}(H/{\mathbb{Q}})\otimes{\mathbb{F}}_{2}$, hence it is onto $\mathrm{Gal}(H/{\mathbb{Q}})/M$. Because $\varepsilon_{f}(c)=(-1)^{k}=-1$, $c,\sigma_{0}$ are two distinct non-trivial $2$-torsion elements in $\mathrm{Gal}(H/{\mathbb{Q}})$, so $(u,v)\in\mathbb{F}_{2}^{\oplus 2}\mapsto\sigma_{0}^{u}c^{v}\in\mathrm{Gal}(H/{\mathbb{Q}})/M$ is onto. Since $\sigma_{0}\notin M$, $\mathrm{Gal}(H/{\mathbb{Q}})$ is then generated by $M$ and $c$ if and only if $\sigma_{0}c\in M$, which is equivalent to the statement that for all $(\gamma,\delta)\in\Gamma_{f}$, $\gamma(i)=\delta(\sigma_{0})\delta(-1)i$.
@@
Suppose finally that $\varepsilon_{f}^{2}=1$. By Lemma 1, for any $(\gamma,\delta)\in\Gamma_{f}$, $\delta$ is a quadratic character, so $\mathrm{Gal}(H/{\mathbb{Q}})$ is a ${\mathbb{F}}_{2}$-vector space; by the previous paragraph, it has dimension two and $(\sigma_{0},c)$ is a basis of this vector space. One can check directly that $\Gamma_{f}$ also has exponent two and that the bilinear pairing $[(\gamma,\delta),\sigma]\in\Gamma_{f}\times\mathrm{Gal}(H/{\mathbb{Q}})\mapsto\delta(\sigma)\in\{\pm 1\}$ is a perfect duality of ${\mathbb{F}}_{2}$-vector spaces. Therefore, the homomorphism $(\gamma,\delta)\in\Gamma_{f}\mapsto\frac{\gamma(i)}{i}\in\{\pm 1\}$ is given by $(\gamma,\delta)\mapsto\delta(\sigma)$ for some $\sigma\in\{\sigma_{0},c,\sigma_{0}c\}$. By taking $(\gamma,\delta)=(c,\varepsilon_{f}^{-1})$, we find that $\varepsilon_{f}(\sigma)=-1$, so $\sigma\in\{c,\sigma_{0}c\}$. Then $i$ is the element $\alpha$ attached to $\sigma$ by Corollary 2, so it follows that $\mathcal{B}(\sigma,\cdot)=1$ i.e. $\sigma\in M$. Because $\sigma_{0}\notin M$, $\sigma$ is the only element of $\{c,\sigma_{0}c\}$ which lies in $M$. Now, by definition, $(d)$ holds if and only if $\sigma=c$, if and only if $\sigma_{0}c\notin M$, if and only if $(e)$ holds.
@@
    Let $f\in\mathcal{S}_{2}(\Gamma_{0}(63))$ be a newform with LMFDB [13] label $63.2.a.b$ [^3] Technically, these labels refer to a Galois orbit of newforms; however, the choice of embedding does not matter in all the situations below, so we omit it.: it has coefficients in ${\mathbb{Q}}(\sqrt{3})$, an inner twist by the quadratic character $\varepsilon$ of $K={\mathbb{Q}}(\sqrt{-3})$ and no complex multiplication. The newform $g\in\mathcal{S}_{1}(\Gamma_{1}(675))$ with LMFDB label $675.1.c.b$ has rational coefficients, character $\varepsilon$ and complex multiplication by $K$. The argument shows that (after possibly removing a finite number of primes) $T_{f,g,\mathfrak{p}}$ does not satisfy (wE) when the residue characteristic is congruent to $5,7$ mod $12$.
@@
    Consider a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(75))$ with LMFDB label $75.3.d.b$. Then $L={\mathbb{Q}}(i,\sqrt{5})$, $F={\mathbb{Q}}$, $H_{0}={\mathbb{Q}}(\sqrt{-15})$, $H={\mathbb{Q}}(j,\sqrt{5})$. Then it can be checked that $\mathrm{Gal}(H/{\mathbb{Q}})$ is generated by $M$ and $c$. When $\varepsilon$ is the character of the unique real quadratic field contained in $H$ and $g$ is a weight one newform with character $\varepsilon\varepsilon_{f}^{-1}$ and with real multiplication by $\varepsilon$, property (sE) then holds for all but finitely many $T_{f,g,\mathfrak{p}}$, even though Theorem Introduction does not apply.
@@
    For a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(24))$ with LMFDB label $24.3.h.c$, the condition $(c)$ of Proposition 3 is satisfied: indeed, in this situation $L={\mathbb{Q}}(\sqrt{2},\sqrt{-7})$ does not contain $i$, while $H={\mathbb{Q}}(\sqrt{2},j)$, and $H_{0}={\mathbb{Q}}(\sqrt{-6})$ is such that $[H:H_{0}]=2$. Indeed, $B$ contains the two distinct non-trivial elements $\mathcal{B}(c,\cdot)$ and $\mathcal{B}(\mathrm{Frob}_{5},\cdot)$ (they are the Kummer characters associated to $\sqrt{14}$ and $\sqrt{2}$ respectively), so that $M$ is trivial and $\mathrm{Gal}(H/{\mathbb{Q}})$ is larger than $\{\mathrm{id},c\}$.
@@
    Even when $i\in L$ is not a value of $\varepsilon_{f}$, it is not true that $(e)$ implies $(d)$. A counter-example is given by the newform $f\in\mathcal{S}_{3}(\Gamma_{1}(324))$ with LMFDB label $324.3.f.s$ [^4] With the tools dicussed here, this verification is not difficult using the LMFDB data and MAGMA [1]; we omit it for the sake of brevity.. In this situation, $\varepsilon_{f}$ has conductor $36$ and order $6$, while $L$ has degree $16$ and $|\Gamma_{f}|=8$, so that $F={\mathbb{Q}}(\sqrt{57})$.
@@
    Consider a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(21))$ with LMFDB label $21.3.h.b$. In this case, one has $|\Gamma_{f}|=4$, but $H=H_{0}={\mathbb{Q}}(j){\mathbb{Q}}(\mu_{7})^{+}$. In fact, $F={\mathbb{Q}}$, $L={\mathbb{Q}}(j,\sqrt{-5})$, and one can directly compute that $B=\{\mathcal{B}(\mathrm{id},\cdot),\mathcal{B}(c,\cdot)\}$.
@@
Let $C$ be the conductor of $\chi$, $D$ be a positive multiple of $C$ and $K^{\prime}/K$ the finite abelian extension whose group of norms is $K^{\times}\prod_{v}{M_{v}}$ where $M_{v}={\mathbb{C}}^{\times}$ if $v$ is complex, ${\mathbb{R}}^{\times}_{+}$ if $v$ is real, $\mathcal{O}_{K_{v}}^{\times}$ if $v$ is finite and not dividing $D$, and $1+D\mathcal{O}_{K_{v}}$ otherwise. Because this group of norms is stable under $\mathrm{Gal}(K/{\mathbb{Q}})$, $K^{\prime}/{\mathbb{Q}}$ is Galois. By functoriality in class field theory [18, Proposition IV.5.9, Sections VI.4, VI.5], $\chi$ vanishes on the kernel of $V_{K^{\prime}}$ as long as the following claim is true: for any $z\in\mathbb{A}_{{\mathbb{Q}}}^{\times}$ whose image in $\mathbb{A}_{K}^{\times}$ is contained in $K^{\times}\prod_{v}{M_{v}}$, we can write $z=\kappa(z_{\infty},z^{\prime})$, where $\kappa\in{\mathbb{Q}}^{\times}$, $(z_{\infty},z^{\prime})\in{\mathbb{R}}^{\times}\otimes(\hat{{\mathbb{Z}}}\cap(1+C\hat{{\mathbb{Z}}}))$, and $z_{\infty}>0$ if $K$ is real.
When $K$ is imaginary, we may assume (after multiplying $z$ with a rational number) that $z=(z_{0},z_{f})$ with $z_{f}\in\hat{{\mathbb{Z}}}^{\times}$ and $z_{0}\in{\mathbb{R}}^{\times}_{+}$, and thus that its image in $\mathbb{A}_{K}^{\times}$ is contained in $\mathcal{O}_{K}^{\times}\prod_{v}{M_{v}}$. It is then enough to find $D$ such that any unit $u\in\mathcal{O}_{K}^{\times}$ which is congruent to an integer modulo $D$ is congruent to ±$\pm 1$ modulo $C$ (this ±$\pm 1$ corresponds to the fact that the complex conjugation is contained in the kernel). Since $\mathcal{O}_{K}^{\times}$ has order $2,4$ or $6$, $D=11C$ works, since ${\mathbb{F}}_{11}$ does not contain primitive third or fourth roots of unity.
When $K$ is real, we may assume (after multiplying $z$ with a rational number) that $z=(z_{0},z_{f})$ with $z_{f}\in\hat{{\mathbb{Z}}}^{\times}$ and $z_{0}\in{\mathbb{R}}^{\times}_{+}$, and thus that its image in $(\mathcal{O}_{K}\otimes\hat{{\mathbb{Z}}})^{\times}$ is congruent to a totally positive unit modulo $D$. It is thus enough to find $D$ such that any totally positive unit $u\in\mathcal{O}_{K}^{\times}$ congruent to an integer modulo $D$ is congruent to $1$ modulo $C$.
Fix a generator of the subgroup $u$ of totally positive units in $\mathcal{O}_{K}$ and suppose that there is no $D$ satisfying the previous condition. Then, with $D_{n}=(n+C)!$, there is an integer $d_{n}$ such that $u^{d_{n}}$ is congruent to an integer modulo $D_{n}$ but not to $1$ modulo $C$. Since $u$ is a totally positive unit, its conjugate under $\mathrm{Gal}(K/{\mathbb{Q}})$ is $u^{-1}$: thus $u^{2d_{n}}$ is congruent to $1$ modulo $D_{n}$. Let $\delta\in\hat{{\mathbb{Z}}}$ be any limit point of the $d_{n}$, then $u^{2\delta}=1$ in $(\mathcal{O}_{K}\otimes\hat{{\mathbb{Z}}})$. By [19, Lemma 8.2.4], one has $02\delta=$ 0 hence $\delta=0$. This implies that, for some $n$, $d_{n}$ is divisible by $|(\mathcal{O}_{K}/C\mathcal{O}_{K})^{\times}|$, so that $u^{d_{n}}$ is congruent to $1$ modulo $C$, a contradiction.
In either case, for a suitable value of $D$, the extension $K^{\prime}$ is such that $\ker{\chi}$ contains $\ker{V_{K^{\prime}}}$. This implies that there is a character $\psi_{0}:\mathrm{Gal}(K^{\prime}/K)\rightarrow{\mathbb{C}}^{\times}$ such that $\psi_{0}\circ V_{K^{\prime}}=\chi$. In particular, if we view $\psi_{0}$ as a character $G_{K}^{ab}\rightarrow\mathrm{Gal}(K^{\prime}/K)\rightarrow{\mathbb{C}}^{\ast}$, one has $\psi_{0}\circ V=\chi$.
@@
Let $S$ be any finite set of finite places of $K$. By Chebotarev, there exists an odd prime $q\equiv-1\pmod{p}$ inert in $K$ such that $q\notin S$. By local class field theory [18, Theorems V.1.3, V.1.4], there exists a surjective character $\theta_{q}:G_{K_{q}}\rightarrow{\mathbb{Z}}/p{\mathbb{Z}}$ whose restriction to the inertia subgroup has order $p$; in particular, if $\theta^{\prime}_{q}$ denotes the conjugate of $\theta$ by the action of $\mathrm{Gal}(K_{q}/{\mathbb{Q}}_{q})$, $\theta^{\prime}_{q}$ and $\theta_{q}^{-1}$ have the same restriction to the inertia group, so that the restriction to the inertia group of $\theta^{\prime}_{q}\theta^{-1}$ has order $p$. By the Grunwald–Wang theorem [19, Theorem 9.2.3], there exists a continuous character $\theta:G_{K}\rightarrow{\mathbb{Z}}/p{\mathbb{Z}}$ extending $\theta_{q}$: in particular, the restriction to the inertia group at $q$ of $\lambda:=e^{\frac{2i\pi}{p}(\theta^{c}-\theta)}$ has order $p$. Moreover, one has $\lambda^{p}=\lambda\lambda^{c}=1$.
@@
Let $i\geq 1$ and $I_{q_{i}}$ be an inertia subgroup fo $G_{K}$ at $q_{i}$. Then $\lambda_{i}(I_{q_{i}})$ is a subgroup of $\lambda_{i}(G_{K})\simeq{\mathbb{Z}}/p{\mathbb{Z}}$ which is not reduced to the unit element, so $\lambda_{i}(I_{q_{i}})=\lambda_{i}(G_{K})$ is the set of $p$-th roots of unity in $K$. Therefore, the character $\alpha_{i}^{-1}\alpha_{i}^{c}=\lambda_{i}^{-2}\psi_{0}^{-1}\psi_{0}^{c}$ is ramified at $q_{i}$. In particular, $\alpha_{i}^{-1}\alpha_{i}^{c}$ is not trivial, so, by Mackey’s criterion [24, Corollary to Proposition 23], $\rho_{i}$ is irreducible.

### v3

@@ 5. $(e)$
Since $H:H_{0}]=2$, the group $\mathrm{Gal}(H/H_{0})$ is cyclic of order two. The group $\mathrm{Gal}(H_{0}/{\mathbb{Q}})$ is isomorphic to the image of $\varepsilon_{f}$ (which contains 1-1), so it is a cyclic group with even cardinality $n$. Fix $\sigma\in\mathrm{Gal}(H/{\mathbb{Q}})$ such that $\varepsilon_{f}(\sigma)$ has order $n$. Then $\sigma^{n}\in\mathrm{Gal}(H/H_{0})\cap M$ because $n$ is even and $\mathrm{Gal}(H/{\mathbb{Q}})/M$ has exponent dividing $2$, so $\sigma^{n}$ is trivial by the second paragraph. Thus $\sigma_{|H_{0}}\mapsto\sigma$ defines a splitting of the above exact sequence, and therefore $\mathrm{Gal}(H/{\mathbb{Q}})={\mathbb{Z}}/2{\mathbb{Z}}\sigma_{0}\oplus{\mathbb{Z}}/n{\mathbb{Z}}\sigma$. Hence, representatives for $\mathrm{Gal}(H/{\mathbb{Q}})/M$ are given (possibly with repetitions) by $\sigma_{0},\sigma_{0}\sigma,\sigma,\mathrm{id}$. If $\sigma$ (or $\sigma\sigma_{0}$) is in $M$, the previous paragraph shows that $\varepsilon_{f}(\sigma)=\varepsilon_{f}(\sigma\sigma_{0})\in L^{\times 2}$. Since $\varepsilon_{f}(\sigma)$ is a generator of the image of $\varepsilon_{f}$, this implies that $\varepsilon_{f}$ takes its values in $L^{\times 2}$. Hence, in the case $(c)$, $M$ does not contain $\sigma$ or $\sigma_{0}\sigma$, so $|B|>2$ and we are done. We assume from now on that $i\in L$.
@@
Note that $(d)$ (resp. $(e)$) holds if and only if $c\in M$ (resp. $\sigma_{0}c\notin M$). Hence, since $\sigma_{0}\notin M$, $(d)$ implies $(e)$. Now, when $(e)$ holds, one has $\sigma_{0}c\notin M$ and $\sigma_{0}\notin M$, so $\sigma_{0}$ is not in the subgroup generated by $c$ and $M$, and the conclusion follows.
@@
If $i$ is not a value taken by $\varepsilon_{f}$ (which is for instance the case if $\varepsilon_{f}^{2}=1$), then $n$ is not divisible by $4$. Because $\mathrm{Gal}(H/{\mathbb{Q}})$ is isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}\oplus{\mathbb{Z}}/n{\mathbb{Z}}$, the $2$-torsion subgroup of $\mathrm{Gal}(H/{\mathbb{Q}})$ is isomorphic to ${\mathbb{F}}_{2}^{\oplus 2}$ and projects isomorphically on $\mathrm{Gal}(H/{\mathbb{Q}})\otimes{\mathbb{F}}_{2}$, hence it is onto $\mathrm{Gal}(H/{\mathbb{Q}})/M$. Because $\varepsilon_{f}(c)=(-1)^{k}=-1$, $c,\sigma_{0}$ are two distinct non-trivial $2$-torsion elements in $\mathrm{Gal}(H/{\mathbb{Q}})$, so $(u,v)\in\mathbb{F}_{2}^{\oplus 2}\mapsto\sigma_{0}^{u}c^{v}\in\mathrm{Gal}(H/{\mathbb{Q}})/M$ is onto. Since $\sigma_{0}\notin M$, $\mathrm{Gal}(H/{\mathbb{Q}})$ is then generated by $M$ and $c$ if and only if $\sigma_{0}c\in M$, if and only if $(e)$ does not hold, and the conclusion follows.
@@
Suppose finally that $\varepsilon_{f}^{2}=1$. By Lemma 1, for any $(\gamma,\delta)\in\Gamma_{f}$, $\delta$ is a quadratic character, so $\mathrm{Gal}(H/{\mathbb{Q}})$ is a ${\mathbb{F}}_{2}$-vector space; by the previous paragraph, it has dimension two and $(\sigma_{0},c)$ is a basis of this vector space. One can check directly that $\Gamma_{f}$ also has exponent two and that the bilinear pairing $(\gamma,\delta),\sigma]\in\Gamma_{f}\times\mathrm{Gal}(H/{\mathbb{Q}})\mapsto\delta(\sigma)\in\{\pm 1\}$ is a perfect duality of ${\mathbb{F}}_{2}$-vector spaces. Therefore, the homomorphism $(\gamma,\delta)\in\Gamma_{f}\mapsto\frac{\gamma(i)}{i}\in\{\pm 1\}$ is given by $(\gamma,\delta)\mapsto\delta(\sigma_{1})$ for some $\sigma_{1}\in\{\sigma_{0},c,\sigma_{0}c\}$. By taking $(\gamma,\delta)=(c,\varepsilon_{f}^{-1})$, we find that $\varepsilon_{f}(\sigma_{1})=-1$, so $\sigma_{1}\in\{c,\sigma_{0}c\}$. Then $i$ is the element $\alpha$ attached to $\sigma_{1}$ by Corollary 2, so it follows that $\mathcal{B}(\sigma_{1},\cdot)=1$ i.e. $\sigma_{1}\in M$. Because $\sigma_{0}\notin M$, $\sigma_{1}$ is the only element of $\{c,\sigma_{0}c\}$ which lies in $M$. Now, by definition, $(d)$ holds if and only if $\sigma_{1}=c$, if and only if $\sigma_{0}c\notin M$, if and only if $(e)$ holds.
@@
    Let $f\in\mathcal{S}_{2}(\Gamma_{0}(63))$ be a newform with LMFDB [14] label 63.2.a.b63.2.a.b [^3] Technically, these labels refer to a Galois orbit of newforms; however, the choice of embedding does not matter in all the situations below, so we omit it.: it has coefficients in ${\mathbb{Q}}(\sqrt{3})$, an inner twist by the quadratic character $\varepsilon$ of $K={\mathbb{Q}}(\sqrt{-3})$ and no complex multiplication. The newform $g\in\mathcal{S}_{1}(\Gamma_{1}(675))$ with LMFDB label $675.1.c.b$ has rational coefficients, character $\varepsilon$ and complex multiplication by $K$. The argument shows that (after possibly removing a finite number of primes) $T_{f,g,\mathfrak{p}}$ does not satisfy (wE) when the residue characteristic is congruent to $5,7$ mod $12$.
@@
    Consider a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(75))$ with LMFDB label $75.3.d.b$. Then $L={\mathbb{Q}}(i,\sqrt{5})$, $F={\mathbb{Q}}$, $H_{0}={\mathbb{Q}}(\sqrt{-15})$, $H={\mathbb{Q}}(\sqrt{-3},\sqrt{5})$. Then it can be checked that $\mathrm{Gal}(H/{\mathbb{Q}})$ is generated by $M$ and $c$. When $\varepsilon$ is the character of the unique real quadratic field contained in $H$ and $g$ is a weight one newform with character $\varepsilon\varepsilon_{f}^{-1}$ and with real multiplication by $\varepsilon$, property (sE) then holds for all but finitely many $T_{f,g,\mathfrak{p}}$, even though Theorem Introduction does not apply.
@@
    For a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(24))$ with LMFDB label $24.3.h.c$, the condition $(c)$ of Proposition 3 is satisfied: indeed, in this situation $L={\mathbb{Q}}(\sqrt{2},\sqrt{-7})$ does not contain $i$, while $H={\mathbb{Q}}(\sqrt{2},\sqrt{-3})$, and $H_{0}={\mathbb{Q}}(\sqrt{-6})$ is such that $H:H_{0}]=2$. Indeed, $B$ contains the two distinct non-trivial elements $\mathcal{B}(c,\cdot)$ and $\mathcal{B}(\mathrm{Frob}_{5},\cdot)$ (they are the Kummer characters associated to $\sqrt{14}$ and $\sqrt{2}$ respectively), so that $M$ is trivial and $\mathrm{Gal}(H/{\mathbb{Q}})$ is larger than $\{\mathrm{id},c\}$.
@@
    Even when $i\in L$ is not a value of $\varepsilon_{f}$, it is not true that $(e)$ implies $(d)$. A counter-example is given by the newform $f\in\mathcal{S}_{3}(\Gamma_{1}(324))$ with LMFDB label 324.3.f.s324.3.f.s [^4] With the tools dicussed here, this verification is not difficult using the LMFDB data and MAGMA [1]; we omit it for the sake of brevity.. In this situation, $\varepsilon_{f}$ has conductor $36$ and order $6$, while $L$ has degree $16$ and $|\Gamma_{f}|=8$, so that $F={\mathbb{Q}}(\sqrt{57})$.
@@
    Consider a newform $f\in\mathcal{S}_{3}(\Gamma_{1}(21))$ with LMFDB label $21.3.h.b$. In this case, one has $|\Gamma_{f}|=4$, but $H=H_{0}={\mathbb{Q}}(\sqrt{-3},\mu_{7})^{+}$. In fact, $F={\mathbb{Q}}$, $L={\mathbb{Q}}(\sqrt{-3},\sqrt{-5})$, and one can directly compute that $B=\{\mathcal{B}(\mathrm{id},\cdot),\mathcal{B}(c,\cdot)\}$.
@@
Let $C$ be the conductor of $\chi$, $D$ be a positive multiple of $C$ and $K^{\prime}/K$ the finite abelian extension whose group of norms is $K^{\times}\prod_{v}{M_{v}}$ where $M_{v}={\mathbb{C}}^{\times}$ if $v$ is complex, $\mathbb{R}$$^{\times}_{+}$ if $v$ is real, $\mathcal{O}_{K_{v}}^{\times}$ if $v$ is finite and not dividing $D$, and $1+D\mathcal{O}_{K_{v}}$ otherwise. Because this group of norms is stable under $\mathrm{Gal}(K/{\mathbb{Q}})$, $K^{\prime}/{\mathbb{Q}}$ is Galois. By functoriality in class field theory [19, Proposition IV.5.9, Sections VI.4, VI.5], $\chi$ vanishes on the kernel of $V_{K^{\prime}}$ as long as the following claim is true: for any $z\in\mathbb{A}_{{\mathbb{Q}}}^{\times}$ whose image in $\mathbb{A}_{K}^{\times}$ is contained in $K^{\times}\prod_{v}{M_{v}}$, we can write $z=\kappa(z_{\infty},z^{\prime})$, where $\kappa\in{\mathbb{Q}}^{\times}$, $)(z_{\infty},z^{\prime})\in{\mathbb{R}}^{\times}\times(\hat{{\mathbb{Z}}}^{\times}\cap(1+C\hat{{\mathbb{Z}}}$)), and $z_{\infty}>0$ if $K$ is real.
When $K$ is imaginary, we may assume (after multiplying $z$ with a rational number) that $z=(z_{0},z_{f})$ with $z_{f}\in\hat{{\mathbb{Z}}}^{\times}$ and $z_{0}\in{\mathbb{R}}^{\times}_{+}$, and thus that its image in $\mathbb{A}_{K}^{\times}$ is contained in $\mathcal{O}_{K}^{\times}\prod_{v}{M_{v}}$. It is then enough to find $D$ such that any unit $u\in\mathcal{O}_{K}^{\times}$ which is congruent to an integer modulo $D$ is congruent to $\pm 1$ modulo $C$ (this $\pm 1$ corresponds to the fact that the complex conjugation is contained in the kernel). Since $\mathcal{O}_{K}^{\times}$ has order $2,4$ or $6$, $D=11C$ works, since ${\mathbb{F}}_{11}$ does not contain primitive third or fourth roots of unity.
[…]
In either case, for a suitable value of $D$, the extension $K^{\prime}$ is such that $\ker{\chi}$ contains $\ker{V_{K^{\prime}}}$. This implies that there is a character $\psi_{0}:\mathrm{Gal}(K^{\prime}/K)\rightarrow{\mathbb{C}}^{\times}$ such that $\psi_{0}\circ V_{K^{\prime}}=\chi$. In particular, if we view $\psi_{0}$ as a character $G_{K}^{ab}\rightarrow\mathrm{Gal}(K^{\prime}/K)\rightarrow{\mathbb{C}}^{\times}$, one has $\psi_{0}\circ V=\chi$.
@@
Let $S$ be any finite set of finite places of $K$. By Chebotarev, there exists an odd prime $q\equiv-1\pmod{p}$ inert in $K$ such that $q\notin S$. By local class field theory [19, Theorems V.1.3, V.1.4], there exists a surjective character $\theta_{q}:G_{K_{q}}\rightarrow{\mathbb{Z}}/p{\mathbb{Z}}$ whose restriction to the inertia subgroup has order $p$; in particular, if $\theta^{\prime}_{q}$ denotes the conjugate of $\theta_{q}$ by the action of $\mathrm{Gal}(K_{q}/{\mathbb{Q}}_{q})$, $\theta^{\prime}_{q}$ and $\theta_{q}$ have the same restriction to the inertia group, so that the restriction to the inertia group of $\theta^{\prime}_{q}-\theta_{q}$ has order $p$. By the Grunwald–Wang theorem [20, Theorem 9.2.3], there exists a continuous character $\theta:G_{K}\rightarrow{\mathbb{Z}}/p{\mathbb{Z}}$ extending $\theta_{q}$: in particular, the restriction to the inertia group at $q$ of $\lambda:=e^{\frac{2i\pi}{p}(\theta^{c}-\theta)}$ has order $p$. Moreover, one has $\lambda^{p}=\lambda\lambda^{c}=1$.
@@
Let $i\geq 1$ and $I_{q_{i}}$ be an inertia subgroup of $G_{K}$ at $q_{i}$. Then $\lambda_{i}(I_{q_{i}})$ is a subgroup of $\lambda_{i}(G_{K})\simeq{\mathbb{Z}}/p{\mathbb{Z}}$ which is not reduced to the unit element, so $\lambda_{i}(I_{q_{i}})=\lambda_{i}(G_{K})$ is the set of $p$-th roots of unity in $\mathbb{C}$}. Therefore, the character $\alpha_{i}^{-1}\alpha_{i}^{c}=\lambda_{i}^{-2}\psi_{0}^{-1}\psi_{0}^{c}$ is ramified at $q_{i}$. In particular, $\alpha_{i}^{-1}\alpha_{i}^{c}$ is not trivial, so, by Mackey’s criterion [25, Corollary to Proposition 23], $\rho_{i}$ is irreducible.